### 30 Jan 2024 v1

# EdgeOL: Efficient in-situ Online Learning on Edge Devices

Abstract.
Emerging applications, such as robot-assisted eldercare and object recognition, generally employ deep learning neural networks (DNNs) models and naturally require: i) handling streaming-in inference requests and ii) adapting to possible deployment scenario changes. Online model fine-tuning is widely adopted to satisfy these needs. However, fine-tuning involves significant energy consumption, making it challenging to deploy on edge devices. In this paper, we propose EdgeOL, an edge online learning framework that optimizes inference accuracy, fine-tuning execution time, and energy efficiency through both inter-tuning and intra-tuning optimizations. Experimental results show that, on average, EdgeOL reduces overall fine-tuning execution time by 82%, energy consumption by 74%, and improves average inference accuracy by 1.70% over the immediate online learning strategy.
11footnotetext: Equal contribution.
1. Introduction
Deep learning neural networks (DNNs) have gained significant popularity in emerging application domains such as robot-assisted eldercare (do2018rish, ; bemelmans2012socially, ), object recognition (doshi2020continual, ; pellegrini2020latent, ), and wild surveillance (akbari2021applications, ; ke2020smart, ). These applications generally deploy DNN models on energy-constrained edge devices, such as mobile handhelds and internet-of-things (IoT) devices (chen2019deep, ; li2018learning, ; wang2018deep, ; deng2014deep, ). There are two fundamental requirements of deploying DNN models on edge devices: (1) adaptiveness and (2) energy-efficiency. From the perspective of adaptiveness, the DNN applications commonly have streaming-in training data and inference requests over time. This requires model fine-tuning using incoming training data to i) adapt to scenario changes, and meanwhile ii) maximize the average inference accuracy for the streaming-in inference requests (details in Section 2). For example, to maintain high accuracy, an object recognition system needs a timely update of its DNN model when working under different environments or conditions, such as involving new classes of data or instances of existing classes but with new patterns (e.g., strong fog or snow) (belgiovine2022hri, ; antoni2023embedded, ; lomonaco2017core50, ; lomonaco2020rehearsal, ), while keeping the recognition functions online. Regarding energy efficiency, the DNN applications need to optimize energy efficiency since they are often deployed on edge devices with constrained power capacities (wang2019e2, ; yang2017designing, ), such as battery-powered camera drones or mobile phones.
Existing approaches usually employ online learning to ensure model adaptiveness. Immediate online learning is an extreme case of online learning that performs immediate model fine-tuning once new training data arrives (nallaperuma2019online, ; gomes2019machine, ; fontenla2013online, ; hayes2020lifelong, ). As a result, it guarantees high inference accuracy for incoming inference requests since the model is always up-to-date. However, this requires a large amount of computation as well as significant memory overheads from frequent model loading, saving, and system initialization, making it less energy efficient. While fine-tuning (i.e., training) models at a fixed and lower frequency seems to be a reasonable trade-off for accuracy and energy efficiency, it is far from an optimal solution (details discussed in Section 3.2).
In this paper, we aim to achieve both adaptiveness and energy efficiency by proposing a framework called EdgeOL. Our design motivation stems from the observation that there are redundant computation and memory accesses during the fine-tuning stage in immediate online learning. Specifically, we first observe that many fine-tuning rounds contribute little to the inference accuracy. That is, selectively delaying and merging some fine-tuning rounds and reducing the fine-tuning frequency will not hurt the inference accuracy. We call this inter-tuning redundancy. Second, we observe that some layers reach convergence during fine-tuning. Thus, freezing those layers will not affect the inference accuracy. We call this intra-tuning redundancy. Moreover, freezing layers helps avoid over-adaptation to training data and improve the model convergence speed by reducing the number of weights being trained. This allows the streaming-in inference requests to use a robust model with a higher inference accuracy while reducing the fine-tuning time and energy.
To summarize, we make the following contributions.
  • $\bullet$
    We conduct comprehensive characterization that quantifies the fine-tuning execution time, energy efficiency, and inference accuracy of existing online learning techniques. We reveal that there exist substantial inter-tuning and intra-tuning redundancies that can be optimized to significantly reduce the fine-tuning execution time and energy consumption while improving the inference accuracy.
  • $\bullet$
    We propose EdgeOL that consists of: i) inter-tuning optimization that dynamically and adaptively determines the fine-tuning frequency, and ii) similarity-guided layer freezing for intra-tuning optimization.
  • $\bullet$
    We evaluate EdgeOL using various DNN models and datasets in both computer vision (CV) and natural language processing (NLP) domains. Experimental results show that, compared to immediate online learning in CV domain, EdgeOL saves 82% (76% in NLP domain) of overall fine-tuning execution time and 74% (63% in NLP domain) of energy consumption on average. Furthermore, EdgeOL improves the average inference accuracy of all streaming-in inference requests by 1.70% (1.76% in NLP domain).
  • $\bullet$
    We make a comprehensive evaluation of EdgeOL under different representative scenario change types in online learning, including task change and input distortion. The results show the effectiveness and great generalizability of EdgeOL for different types of application scenarios.
  • $\bullet$
    We demonstrate EdgeOL outperforms state-of-the-art efficient training methods, including layer freezing frameworks i) Egeria and ii) AutoFreeze, iii) sparse training framework RigL, and iv) efficient online learning framework Ekya, even if they have been optimized by our proposed inter-tuning optimization. EdgeOL provides 2.1$\times$, 2.2$\times$, 2.6$\times$, and 2.2$\times$ energy savings, respectively, while delivering 1.76%, 2.10%, 2.17%, and 1.89% higher average inference accuracy.
2. Background
Online learning. As shown in Figure 1, online learning can be categorized into immediate online learning and fixed-frequency online learning that initiates fine-tuning process after receiving a certain fine-tuning data batches (e.g., two batches in the example in Figure 1). The example involves eight received fine-tuning data batches represented by eight red lines, respectively. The green lines indicate four inference requests. Note that, in practice, inference requests might arrive in bursts (e.g., at $t_{3}$ and $t_{4}$). In immediate online learning, model fine-tuning is triggered once fine-tuning data (i.e., training data) arrives. Thus, the model is fine-tuned eight times in this example. On the other hand, fixed-frequency online learning fine-tunes the model four times in this case. In general, immediate online learning achieves the highest average inference accuracy by frequently updating the model. However, this involves significant overheads from frequent model loading, saving, and system initialization, making it less energy efficient. While fixed-frequency fine-tuning seems to be a reasonable tradeoff between accuracy and energy efficiency, it lacks flexibility and adaptiveness to different cases and is far from an optimal solution (details discussed in Section 3.2).
It is worth mentioning that, for edge online learning systems that employ a supervised learning paradigm, there are several different methods to address the data labeling issue for the newly arrived training data. For example, some systems label the training data using a highly accurate but expensive model (with deeper architecture and a larger size) (bhardwaj2022ekya, ; khani2023recl, ; khani2021real, ; Mullapudi_2019_ICCV, ), and this is essentially that of supervising a low-cost “student” model with a high-cost “teacher” model (knowledge distillation) (gou2021knowledge, ; cho2019efficacy, ; shmelkov2017incremental, ). The reason why we need to train a small model is that the large model cannot keep up with inference on the edge. Moreover, the training data can also be labeled by open-source labeling platforms (soghoyan2021toolbox, ; chew2019smart, ; fiedler2019imagetagger, ) or stand-alone labeling service providers (humansignal2023, ; telusinternational2023, ; imerit2023, ).
Average inference accuracy. In online learning, the ongoing model fine-tuning and continuous arrival of inference requests necessitate an evaluation metric to assess the effectiveness of fine-tuning during the entire online learning process. Thus, the average inference accuracy, which is the arithmetic mean of (instantaneous) inference accuracies for all requests, is commonly used to serve the purpose (bhardwaj2022ekya, ). For example, inference requests occur at times $t_{3}$, $t_{4}$, $t_{6}$, and $t_{11}$, with corresponding accuracies $A_{t3}$, $A_{t4}$, $A_{t6}$, and $A_{t_{11}}$. The average accuracy is thus calculated as $(A_{t3}+A_{t4}+A_{t6}+A_{t{11}})/4$.
Figure 1. Examples of immediate online learning and fixed fine-tuning frequency online learning.
Figure 2. Computation of DNN training.
[…]
Scenario change. We categorize the scenario changes into two types: task change and input distortion. For the task change, the initial model is originally trained on a large and general dataset (e.g., ImageNet), then deployed to work on a new dataset (e.g., CIFAR-10) that could contain the data from the classes that are not observed during the previous training. The task change usually happens when a general model needs to be deployed to a more task-specific application (weiss2016survey, ). The input distortion occurs when a model is trained on a particular task/dataset but needs to handle the inference data that are severely distorted, leading to a significant distribution shift (hendrycks2019benchmarking, ). For example, input data for an object recognition system is strongly distorted by natural environments, such as strong fog or snow, which are not considered in the original model training. Conducting online fine-tuning with the new dataset or using the distorted data can effectively mitigate the effect of scenario change and improve accuracy. In our work, we comprehensively evaluate our method for both types of scenario changes.
Figure 3. Overall fine-tuning execution time and energy consumption of nine strategies.
Figure 4. Average inference accuracy using nine strategies.
Figure 3. Overall fine-tuning execution time and energy consumption of nine strategies.
Figure 4. Average inference accuracy using nine strategies.
Figure 5. Accuracy improvement curve of ResNet50 and MobileNetV2 using strategy $S7$.
3. Challenges and Opportunities
First, we quantitatively characterize the impact of different online learning strategies on the fine-tuning execution time, energy consumption, and inference accuracy. The differences between these learning strategies primarily target two aspects: 1) inter-tuning and 2) intra-tuning.
3.1. Experimental Setup
Platform: We use the NVIDIA Jetson Xavier NX and choose the 15W 6-Core power mode with max GPU speed.
Model and dataset: In this section, we employ two popular DNN models, ResNet50 (Res50 for short) (he2016deep, ) and MobileNetV2 (MBV2 for short) (sandler2018mobilenetv2, ). For clarity and simplicity, we use “ImgNet to CF10” as an example test case of scenario change in this section. Specifically, the initial model was well-trained on the ImageNet dataset (deng2009imagenet, ) and will be online fine-tuned and evaluated on CIFAR-10 dataset (krizhevsky2009learning, ) with streaming-in training data (from the training set, containing 50,000 images in total) and inference requests (from the testing set). More experiments are included in Section 5.4 to cover more complex test cases like CORe50 benchmarks (lomonaco2020rehearsal, ), which involve multiple scenario changes. The training batch size is fixed to 8 to avoid out-of-memory errors.
Metrics: We use three metrics for evaluation: overall fine-tuning execution time, overall energy consumption, and average inference accuracy. The overall fine-tuning execution time and energy consumption refer to the total time and energy costs during the entire online learning process. They sum up the fine-tuning execution time and energy consumption of all fine-tuning rounds. The average inference accuracy is the average of accuracies over all inference requests.
Fine-tuning Setting: The arrival granularity of training data is 1 batch each time, and we assume there are a total of 500 inference requests. The arrival rate for both the training data and inference requests follows a Poisson distribution to mimic real application scenarios (mattson2020mlperf, ). We also study different numbers of inference requests and different arrival distributions in Section 5.3. We follow the setting in (lomonaco2017core50, ; lomonaco2020rehearsal, ), where each training sample is used only once in the experiment.
Table 1. The configurations of the fine-tuning strategies.
Strategy
Immed.
S1
S2
S3
S4
S5
S6
S7
S8
Number of batches to trigger a fine-tuning round
1
5
10
20
50
100
200
500
1,000
Number of fine-tuning rounds triggered
6,250
1,250
625
313
125
63
32
13
7
3.2. Inter-tuning
We first investigate whether each round of fine-tuning uniformly impacts inference accuracy. As both training data and inference requests are arriving over time, we argue that each fine-tuning round does not contribute equally. As such, those fine-tuning rounds that have little impact on accuracy can be delayed and merged to save time and energy. However, it is non-trivial to determine which rounds are appropriate to be delayed. To this end, we conduct a quantitative study by varying the fine-tuning frequency to reveal these opportunities. Note, the total amount of training data used is not changed as we only delay and merge some fine-tuning rounds. We do not skip using any training data.
We define the fine-tuning frequency as the number of the triggered fine-tuning rounds during a certain period. We only trigger a fine-tuning round after a fixed number of training data batches arrive. With higher fine-tuning frequency, fewer batches are required to trigger a fine-tuning round. As shown in Table 1, we choose nine different fine-tuning frequencies to perform a sweep from immediate online learning (Immed.) to the less frequent fine-tuning strategies $S8$.
3.2.1. Overall fine-tuning execution time and energy consumption
We investigate the overall fine-tuning execution time and energy consumption of the nine strategies during the entire online learning process. All the results are normalized to Immed. As shown in Figure 5, the execution time can be broken down into four parts: 1) system initialization (e.g., model compilation), 2) model loading & saving, 3) data loading, and 4) forward and backward propagation (FP+BP). One can observe that the forward and backward propagation time (grey bar), and data loading time (red bar) are similar across different fine-tuning strategies. This is because the training batch size and the number of total training samples are the same. However, the system initialization time (orange bar), model loading & saving time (blue bar)—the execution overheads that come with each fine-tuning round—grow significantly with increasing fine-tuning frequency. These execution overheads can even dominate the overall execution time and lead to a considerable increase in energy consumption. For example, the energy consumption increases significantly when the fine-tuning frequency increases from $S4$ to Immed. Besides, these overheads vary between models—they account for a higher proportion of the total time in ResNet50 than in MobileNetV2 because ResNet50 (100MB) is 7$\times$ larger than MobileNetV2 (14MB). Additionally, reducing fine-tuning frequency can also effectively save execution time and energy consumption.
3.2.2. Average inference accuracy
While less frequent fine-tuning reduces execution time and energy consumption, it could severely degrade accuracy, which is unacceptable, as shown in Figure 5. Using a higher fine-tuning frequency can help improve the accuracy, but there are diminishing returns as it approaches Immed. The accuracy effectively saturates at a relatively high fine-tuning frequency (e.g., Immed. $\sim S2$ for ResNet50 and Immed. $\sim S1$ for MobileNetV2 in Figure 5). The saturation frequency varies between models due to different convergence speeds and upper-bound accuracies.
3.2.3. Accuracy contribution of each fine-tuning round
Figure 5 displays the model validation accuracy (details of validation accuracy are in the Section 4.1.2) trend over all fine-tuning rounds using strategy $S7$ (other strategies showed similar trends) where fine-tuning is triggered if 500 new training batches arrive (see Table 1). The accuracy improves quickly in early fine-tuning rounds, and the improvement slows in later rounds as the models reach a steady state where the accuracy saturates. This demonstrates the potential to decrease the fine-tuning frequency as the model becomes well-tuned. For instance, some rounds can be delayed and merged with marginal accuracy loss in the steady state. However, it should be noted that the number of fine-tuning rounds where accuracy saturates varies across models, as shown in Figure 5, implying the need for adaptive fine-tuning strategies.
3.3. Intra-tuning
For a given round of fine-tuning, we investigate if we can further reduce computation by studying different layers’ impacts on accuracy. A recent study has revealed that some DNN layers show a higher representational similarity even if the models are trained on different datasets (kornblith2019similarity, ). Inspired by this, we explore if we can reduce computation costs without compromising accuracy by “freezing” some layers.
A Preliminary Exploration. We conduct a preliminary experiment to show the effect of layer freezing on execution time, energy consumption, and accuracy. As an example, we employ ResNet50 and MobileNetV2 on the “ImgNet to CF10” test case and follow the experimental setup in Section 3.1, and we use the fine-tuning strategy $S7$.
Figure 6. Fine-tuning time and energy when freezing different numbers of layers.
First, we randomly freeze 10, 30, and 50 CONV layers (with the corresponding BN layers) throughout the entire online learning process.
Figure 6 depicts the execution time and energy consumption when freezing different numbers of layers of ResNet50. We normalized the results to a non-freezing baseline. One can observe that layer freezing effectively reduces the time and energy of the online learning process and the savings increase as more layers are frozen. Importantly, these time and energy savings can be directly achieved using native DNN training frameworks (e.g., PyTorch) and do not require any support of dedicated libraries (e.g., sparse computation) or specific hardware accelerators.
Figure 7. Average inference accuracy when freezing different numbers of layers.
Figure 7 shows the accuracy when freezing different numbers of layers of ResNet50 and MobileNetV2. For a more comprehensive exploration, we randomly selected two groups of layers to freeze, denoted as Config 1 and Config 2, respectively. First, we observe that as more layers are frozen, the accuracy drops accordingly. However, when we compare the accuracy of the two configurations, even with the same number of frozen layers, there is still a significant difference in accuracy. The most intriguing observation is that, when freezing layers appropriately (e.g., freezing 10 layers), the accuracy is not decreased and, in fact, increases over the non-freezing baseline. This is because freezing layers can i) avoid over-adaptation to training data and ii) improve model convergence speed (details in Figure 14 and Section 5.1). This indicates that layer freezing is not a simple trade-off between accuracy and efficiency. As the preliminary experimental results, we use relatively simple settings to explore the feasibility and potential of incorporating layer freezing in online learning. Note that similar trends are observed with other experimental settings (e.g., using other fine-tuning strategies).
Even though layer freezing has promising performance in time and energy efficiency and accuracy in online learning, several critical questions remain. Firstly, to incorporate layer freezing in the fine-tuning process, determining which layers are appropriate to freeze is a challenging problem. Secondly, there are also many choices about when to freeze a layer.
4. EdgeOL Design
Table 2. Abbreviation Description.
Abbreviation
Description
batches_ava
Number of data batches available for fine-tuning
batches_needed
Number of data batches needed to trigger a fine-tuning round
acc_im_this_ft
Validation acc. improvement in this fine-tuning round
acc_im_last_ft
Validation acc. improvement in last fine-tuning round
acc_im_next_ft
Approx. validation acc. improvement in next fine-tuning round
freeze_interval
The interval (iterations) to conduct a freezing process
CKA_variation
The variation rate of CKA
CKA_TH
CKA variation rate threshold to regard CKA is stable
Based on these insights, we propose EdgeOL, an efficient online learning framework for edge devices. Figure 8 shows the overview of the EdgeOL framework to achieve energy efficiency and high inference accuracy through i) inter-tuning and ii) intra-tuning optimizations. Specifically, for inter-tuning, we propose a novel Dynamic and Adaptive Fine-tuning Frequency (DAF) design that dynamically and adaptively adjusts the fine-tuning frequency to reduce the execution time and energy consumption (Section 4.1). For intra-tuning, we propose a similarity-guided freezing (SimFreeze) method to automatically freeze layers during online learning to save computation and memory costs while improving accuracy (Section 4.2). If there exist multiple changes in the deployment scenario during online learning, EdgeOL can also detect them and quickly adapt to the new scenarios (Section 4.3). Moreover, it can utilize unlabeled data via semi-supervised learning (Section 4.4). The EdgeOL optimization design is described in Algorithm 1 with terminology and abbreviations listed in Table 2.
Figure 8. Overview of EdgeOL. ①, ②, and ③ in Figure 8b indicate the occurrence of freezing, matching the case 1, 2, and 3 in Figure 2, respectively. ④ indicates the occurrence of unfreezing right after scenario change is detected.
4.1. Dynamic and Adaptive Fine-tuning Frequency (DAF)
4.1.1. Key Design Factors
Cost-Effectiveness. As discussed in Section 3.2 and Figure 5, each time launching a fine-tuning round will inevitably introduce extra time and energy overheads. Therefore, we must consider the cost-effectiveness. It indicates whether the potential model accuracy improvement of launching a fine-tuning round is worth the incurred overheads.
Inference Intensity. Inference intensity is also closely related to the performance of online learning in real-world applications. Since each incoming inference request is served using the model at that time with instantaneous model accuracy, it would be more desirable to keep the model up-to-date when the inference requests are frequent. Therefore, we take the “inference intensity” into account for adaptive fine-tuning frequency adjustment, improving the practicality of our method.
4.1.2. DAF Design Principle
To consider both cost-effectiveness and inference intensity, the proposed DAF dynamically orchestrates the fine-tuning frequency based on training data availability, the trend of the model’s validation accuracy, and the intensity of inference requests. It is important to emphasize that validation accuracy differs from inference accuracy of inference requests. Validation accuracy is obtained by evaluating model on validation dataset, which is formed by randomly separating a small portion ($\sim$5%) of the streaming-in training data while maintaining the same data distribution (ponce2006dataset, ). We cannot use inference accuracy because, in real-world applications, the inference requests will not have the corresponding ground truth labels; thus, we use validation accuracy to indicate model performance.
Specifically, DAF controls the fine-tuning frequency by using a tunable variable $batches\_needed$. A fine-tuning round is triggered only if the available streaming-in training data reaches $batches\_needed$ (line 2 in Algorithm 1). A larger $batches\_needed$ indicates a lower fine-tuning frequency, wh- ere more fine-tuning rounds are delayed and merged. In our design, the initial value of $batches\_needed$ is the same as immediate online learning (i.e., 1 batch). And we use the following principles to adaptively tune up/down the $batches\_needed$ during online learning.
Tuning down the fine-tuning frequency. In general, as the model gradually converges through multiple fine-tuning rounds during online learning, the cost-effectiveness of the fine-tuning rounds decreases (see Section 3.2.3). Therefore, the DAF design monitors the improvements in the validation accuracy of each fine-tuning round. If the improvement is insufficient to remain cost-effective, the fine-tuning frequency is tuned down accordingly (lines 12 to 17 in Algorithm 1).
Tuning up the fine-tuning frequency. To effectively improve inference accuracy under an intensive inference period, we tune up the fine-tuning frequency in DAF via a popular logarithmic-based function (manaseer2009logarithmic, ). It is calculated as $d=d*(1-1/log(d))$, where $d$ represents the number of data batches needed to trigger a fine-tuning. If inference requests are intensive, the $batches\_needed$ will be decreased, thereby tuning up the fine-tuning frequency (lines 20 to 23 in Algorithm 1). We opt for a logarithmic-based function to decrease $batches\_needed$ because it provides a moderate adjustment compared to two other prevalent value-adjusting functions: exponential-based function (kwak2005performance, ) and additive-based function (dumas2002markovian, ). The logarithmic-based function is less aggressive than the exponential-based function, yet more aggressive than the additive-based function.
Figure 9. An example of adaptive adjustment using DAF.
4.1.3. Case study
Figure 9 shows a real example of how DAF adaptively adjusts the fine-tuning frequency. The results are obtained on ResNet50 with a “ImgNet to CF10” task change scenario. From the figure, we have the following observations: ① shows that $batches\_needed$ has small values and remains unchanged for several consecutive fine-tuning rounds. This is because significant accuracy improvements are achieved at the beginning of the online learning process, and our DAF infers to keep a high fine-tuning frequency. ② shows obvious decreases in $batches\_needed$ as the DAF responds to the intensive incoming inference requests at those moments. Moreover, DAF increases the fine-tuning frequency to keep the model up-to-date. ③ shows the overall increasing trend of $batches\_needed$ throughout the online learning process since the model has generally converged, and DAF decreases the fine-tuning frequency (increases $batches\_needed$) to facilitate higher cost-effectiveness and reduce energy consumption.
1 # Fine-tuning if $batches\_ava\geq batches\_needed$ then
2       trigger_fine_tuning();
3       # SimFreeze for every $freeze\_interval$ training iterations do
4             for each active layer do
5                   cka_calculation();
6                   if $CKA\_variation\leq CKA\_TH$ then
7                         freeze_layer();
8
9
10            $freeze\_interval\leftarrow freeze\_interval\times(1-1/log(freeze\_interval))$ ;
11
12      # Tuning down the fine-tuning frequency if fine-tuning ends then
13             if $acc\_im\_this\_ft\leq 0$ $\mathrm{or}$ $acc\_im\_last\_ft\leq 0$ then
14                   $batches\_needed\leftarrow batches\_needed\times 2$;
15
16            else if $acc\_im\_this\_ft\leq acc\_im\_last\_ft$ then
17                   $acc\_im\_next\_ft\leftarrow acc\_im\_this\_ft\times(acc\_im\_this\_ft/acc\_im\_last\_ft)$;
18                   $batches\_needed\leftarrow batches\_needed\times(acc\_im\_this\_ft/acc\_im\_next\_ft)$;
19
20
21
22# Inference # Tuning up the fine-tuning frequency if inference arrives then
23       do_inference();
24       if inference ends then
25             $batches\_needed\leftarrow batches\_needed\times(1-1/log(batches\_needed))$ ;
26
27
28# Multiple scenario changes if a scenario change is detected then
29       reset_fine-tuning_frequency();
30       update_CKA_test_data();
31       for each frozen layer do
32             comp_cka_with_previous&new_scenario_data();
33             if $CKA\_variation\geq CKA\_TH$ then
34                   unfreeze_layer();
35
36
37
Algorithm 1 EdgeOL
4.2. Similarity-Guided Freezing (SimFreeze)
We next design SimFreeze that adaptively freezes appropriate layers during online learning.
Utilizing self-representational similarity to guide freezing. SimFreeze uses a layer’s self-representational similarity to guide whether it can be frozen. We consider the initial model before fine-tuning as the reference model. We define the self-representational similarity of a layer as the degree of similarity between the output feature maps of a layer in the current model version and the output feature maps of that layer in the reference model. As fine-tuning proceeds, the layers of the model are updated over time, and their self-representational similarity is also recorded. When a layer’s self-representational similarity is stabilized, then we consider the layer to have converged and it can be frozen.
To measure the self-representational similarity of two layers from two models, we use a recently proposed metric Centered Kernel Alignment (CKA) (kornblith2019similarity, ). The CKA value is obtained by comparing the output feature maps of two layers using the same input image batch. It can be calculated as:
(1) $CKA\left(X,Y\right)=\left\|Y^{\mathrm{T}}X\right\|_{\mathrm{F}}^{2}/\left(\left\|X^{\mathrm{T}}X\right\|_{\mathrm{F}}\left\|Y^{\mathrm{T}}Y\right\|_{\mathrm{F}}\right)$
where $X$ and $Y$ are the output feature maps from two layers, and $\left\|\cdot\right\|_{\mathrm{F}}^{2}$ represents the square of the Frobenius norm. A higher CKA value represents that the two layers can generate more similar output feature maps using the same inputs. Moreover, if the CKA value of a layer stabilizes as fine-tuning proceeds, we consider this layer to have converged.
Since calculating CKA requires using the same input data, we collect the first arrived training data batch as our CKA test data. The CKA test data will be used as the input for the models to generate output feature maps of each layer. As shown in Algorithm 1 (lines 5 to 7), periodically (e.g., every 300 iterations), we calculate the CKA and check the self-representational similarity for each active (non-frozen) layer, which is the first step of the freezing process in Figure 8b. For each active layer, we compare the CKA value of the current model to the CKA value calculated last time. The layers whose CKA variations are below the threshold (e.g., 1%) are considered converged and will be frozen (lines 8 and 9 in Algorithm 1), as the second freezing step in Figure 8b.
Figure 10 shows an example of the trend of CKA values of layers 1, 10, 15, 40, and 50 as fine-tuning proceeds. It can be observed that different layers require a different number of training iterations to converge. For example, layer 1 converges at the very beginning, while layer 50 continues to fluctuate at the end of the fine-tuning. Moreover, it is interesting to observe that later layers can converge faster than earlier layers (e.g., layer 15 vs. layer 10). This is due to residual connections in the model network architecture, making some later layers behave like earlier layers (veit2016residual, ). These observations show the feasibility and necessity of freezing layers in an adaptive manner rather than sequentially from front to back. Besides, we can also observe that once a layer has converged, its CKA value will remain stable and will not fluctuate significantly again. Therefore, if a layer is frozen, then it would be good to keep it frozen for higher energy efficiency unless there are changes in the model deployment scenario (discussed in Section 4.3).
Figure 10. CKA variation curve as fine-tuning proceeds. Result is obtained by fine-tuning ResNet50 on “ImgNet to CF10” test case.
Dynamic interval to conduct freezing processes. SimFreeze continuously conducts freezing processes (i.e., tracks the CKA variation and freezes the converged layers). However, conducting the freezing process too frequently (e.g., every training iteration) can lead to excessive overhead, as CKA calculation requires time and energy. Therefore, we opt for periodically tracking the CKA variation of active layers and freezing the converging ones. Further, we propose using a logarithmic-based function that progressively decreases the interval before subsequent freezing process after each freezing process (line 10 in Algorithm 1). This approach takes into consideration that more and more layers converge as fine-tuning proceeds. For example, layers 10 and 40 in Figure 10 have not converged at early training iterations but gradually converged. The log-based function is calculated as $n=n*(1-1/log(n))$, where $n$ represents the interval to conduct a freezing process.
4.3. Handling Multiple Scenario Changes
It is possible that the model undergoes multiple deployment scenario changes during online learning. It is crucial for the model to be updated quickly to adapt to these changes to deliver satisfactory results. In general, the scenario changes can be identified using many different methods. For example, one can track the distribution difference of the data in streaming-in inference requests using methods such as Least-Squares Density Difference (LSDD) (gama2014survey, ) or Maximum Mean Discrepancy (MMD) (gretton2012kernel, ). Or it can also be detected by a stand-alone sensor module in a comprehensive system (e.g., autonomous vehicle system (gormer2009vision, )). EdgeOL is compatible with these detection methods, and it is not our focus in this work.
Once the scenario change is detected, at the inter-tuning level, we reset the fine-tuning frequency to the initial default value, thereby increasing the fine-tuning frequency. At the intra-tuning level, it may not be necessary to unfreeze all the frozen layers since some front layers are task-agnostic (discussed in Section 3.3). To decide which layers to unfreeze, we first update the CKA test data with the new scenario data (line 27 in Algorithm 1). Then, for each frozen layer, we compute its CKA using both original and new scenario CKA test data. If the CKA variation of a layer exceeds the threshold after the scenario change, it indicates that the feature extraction ability of that layer for new scenario data and previous scenario data is significantly different. In this case, we unfreeze that unstable layer to allow it to adapt to the new scenario (lines 28 to 31 in Algorithm 1).
Figure 11. Overall fine-tuning execution time.
Figure 12. Overall fine-tuning energy consumption.
4.4. Utilizing Unlabeled Data
It is possible that some streaming-in training data is not labeled. To fine-tune the model in this case, EdgeOL applies a semi-supervised learning technique (chapelle2009semi, ). In each fine-tuning round, EdgeOL first fine-tunes the model using unlabeled data via the self-supervised learning (chen2020simple, ) to improve the feature extraction ability of the model. Then, EdgeOL fine-tunes the model by supervised learning using the labeled data to improve the model’s performance in the target task (e.g., image classification for a particular dataset).
5. Evaluation
In this section, we evaluate the proposed EdgeOL framework using popular online learning workloads from both CV and NLP domains. The evaluation is conducted under two scenario change types: task change and input distortion.
In CV domain, we employ two popular CNN models ResNet-50 (Res50 for short) and MobileNetV2 (MBV2 for short), as well as a vision transformer model DeiT (tiny version) (pmlr-v139-touvron21a, ). To evaluate EdgeOL on task change, we use two test cases “ImgNet to CF10” and “ImgNet to CalT101”. The models were originally well-trained on the ImageNet dataset and then targeted to serve CIFAR-10/Caltech101 dataset with the corresponding streaming-in training data and inference requests. To evaluate EdgeOL on input distortion, we use the ImageNet-C/CIFAR-10-C dataset (hendrycks2019benchmarking, ), which has the same image classes as ImageNet/CIFAR-10, but its image involves distortion such as fog and snow. Two test cases “ImgNet to ImgNet-C” and “CF10 to CF10-C” are used in our evaluation. We also evaluate EdgeOL on CORe50 benchmarks (lomonaco2020rehearsal, ) for multiple scenario change cases in Section 5.4, which include both task changes and input distortions. In NLP, we evaluate EdgeOL by fine-tuning the BERT-base model (kenton2019bert, ) (originally well-trained on English Wikipedia and the Brown Corpus) using two popular datasets, MRPC (dolan2005automatically, ) and CoLA (warstadt2019neural, ).
Each dataset contains training and testing sets, and a portion of training data (5%) is randomly separated to form a validation dataset as discussed in Section 4.1.2. The experimental setup is the same as in Section 3.1. We assume a total of 500 inference requests, and we provide a sensitivity study on different numbers of inference requests in Section 5.3. All reported accuracies are the average over 5 runs using different random seeds. Unless otherwise stated, the accuracy results refer to the average inference accuracy.
We use immediate online learning as our baseline, where the models are fine-tuned once training data is available. To showcase the benefits of fine-tuning, we present the average inference accuracy achieved by using the pre-trained model without fine-tuning (denoted as “w/o FT”) to serve all the inference requests. We also compare EdgeOL with state-of-the-art efficient training methods (Section 5.2), including layer freezing methods i) Egeria (wang2023egeria, ) and ii) AutoFreeze (liu2021autofreeze, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ).
5.1. Main Results
5.1.1. CV Tasks
Figure 1112, and Table 3 show the overall execution time, energy consumption, and average inference accuracy of the immediate online learning, model without fine-tuning (accuracy only), and our proposed frameworks in CV domain. The execution time and energy consumption are normalized to $Immed.$ As shown in Table 3, fine-tuning can uniformly outperform the model without fine-tuning in accuracy, as it can ensure adaptiveness for the model.
Table 3. Average inference accuracy of all methods.
Model
Method
Dataset
ImgNet to ImgNet to ImgNet to CF10 to
CF10 CalT101 ImgNet-C CF10-C
Res50
w/o FT
12.03
2.10
58.23
85.72
Immed.
85.86
66.73
72.95
93.24
DAF
85.73
66.51
72.82
93.21
SimFreeze
88.21
68.92
74.87
95.05
EdgeOL
88.02
68.68
74.69
94.93
MBV2
w/o FT
12.77
1.93
53.45
79.06
Immed.
83.83
61.52
68.34
89.42
DAF
83.62
61.31
68.20
89.05
SimFreeze
85.88
63.59
70.40
91.03
EdgeOL
85.51
63.34
70.23
90.81
DeiT
w/o FT
11.93
2.26
54.61
84.06
Immed.
86.26
71.61
70.19
92.53
DAF
86.29
71.36
70.05
92.37
SimFreeze
88.06
73.48
71.98
93.97
EdgeOL
87.79
73.27
71.77
93.82
Table 4. Computation cost of entire online learning process. Test case:“ImgNet to CF10”. Method Computation (TFLOPs) Res50 MBV2 Immed. 1,236 96 EdgeOL 767 58
[Uncaptioned image] Figure 13. Computation cost for 1 iteration and memory usage as fine-tuning proceeds.
[Uncaptioned image] Figure 14. Convergence of ResNet50 model on “ImgNet to CF10” test case.
DAF. As shown in Figures 11 and 12, DAF saves average 74%, peak 86% execution time, and average 54%, peak 77% energy compared to $Immed$. These savings come from merging and delaying certain fine-tuning rounds, which can effectively reduce the execution overheads (by 95% on average), including system initialization and model loading & saving (orange bar and blue bar in Figure 11). As shown in Table 3, despite the impressive gains in time and energy efficiency, DAF only incurs a minor 0.16% accuracy drop compared to $Immed$. This is because it adaptively determines the fine-tuning frequency that fits the current scenario.
SimFreeze. SimFreeze reduces average 8%, peak 15% execution time and saves average 20%, peak 33% energy compared to $Immed.$, as shown in Figures 11 and 12. These gains stem from the 40% average savings in model computation (i.e., forward and backward propagation) through layer freezing. Notably, SimFreeze also delivers significantly higher accuracy, a 1.91% average increase over $Immed.$, as shown in Table 3. The reasons are two-fold: First, SimFreeze accelerates model convergence (shown in Figure 14) as freezing layers reduce the number of model weights being trained. Second, SimFreeze avoids excessive adaptation to training data by freezing well-trained layers.
EdgeOL. EdgeOL combines DAF and SimFreeze. From Figures 11, 12, and Table 3, compared to $Immed.$, EdgeOL saves average 82%, peak 91% execution time and average 74%, peak 89% energy, and improves accuracy by an average of 1.70%. Note that EdgeOL shows more time (average 4%) and energy (average 10%) savings in the input distortion test cases, as their scenario changes are less pronounced, allowing greater optimization potential in both inter- and intra-tuning.
Computation Cost and Memory Usage. Table 5.1.1 shows the computation cost reduction. Figure 13 (left) shows the computing cost per iteration as the fine-tuning proceeds. Note, computation cost reduction comes from SimFreeze, as DAF only delays and merges fine-tuning rounds. EdgeOL also saves memory since freezing layers can reduce the intermediate data generated during the computation. As shown in Figure 13 (right), EdgeOL can reduce the memory usage by 40% for ResNet50 and MobileNetV2.
Model Convergence Speed. Figure 14 plots the model convergence during entire online learning. We observe that our EdgeOL helps the model converge faster, leading to a higher accuracy compared to immediate online learning.
Overheads. The major overhead of EdgeOL is the CKA calculation in SimFreeze. This overhead is introduced by i) a forward propagation using a batch of data to get the output feature maps, ii) CKA calculation for active layers using the obtained output feature maps. Fortunately, many layers will be frozen as training proceeds, so the computation of CKA decreases over time. Specifically, in our evaluation, SimFreeze incurs $<$3% additional energy for CKA computation, which is negligible compared to 74% energy benefit from EdgeOL. The reported results include these overheads.
5.1.2. NLP Tasks
We evaluate the EdgeOL framework on NLP tasks to showcase its generalizability. As shown in Table 5, EdgeOL achieves over 30% higher accuracy compared to the model without fine-tuning. When compared to immediate online learning, EdgeOL offers an average reduction of 76% in execution time and 63% in energy consumption, while increasing the accuracy by 1.76% on average. These results demonstrate the generalizability and superiority of EdgeOL.
Table 5. Experimental results in NLP workloads.
Method
MRPC CoLA
Acc.
Time
Energy
Acc.
Time
Energy
w/o FT
31.61
N/A
N/A
30.87
N/A
N/A
Immed.
72.90
3,866
11.50
49.02
9,101
23.88
DAF
72.71
1,273
5.48
48.89
2,749
11.10
SimFreeze
74.52
3,459
9.93
51.16
8,573
21.22
EdgeOL
74.38
873
4.11
51.05
2,278
9.26
Figure 15. Results under the different number of inference requests.
Figure 16. Results under different arrival distributions.
Figure 15. Results under the different number of inference requests.
Figure 16. Results under different arrival distributions.
Figure 17. Results under different CKA variation threshold.
5.2. Comparison with State-of-the-art Efficient Learning Methods
We compare EdgeOL with state-of-the-art efficient training methods, including layer freezing methods i) Egeria (wang2023egeria, ) and ii) AutoFreeze (liu2021autofreeze, ), iii) sparse training framework RigL (evci2020rigging, ), and iv) efficient online learning framework Ekya (bhardwaj2022ekya, ). Results are presented in Table 6. All these methods do not consider optimizations of inter-tuning, which significantly limits their benefits in execution time, energy consumption, and accuracy. For a thorough comparison, we integrate our inter-tuning optimization, DAF, into all methods with identical configurations. Table 6 shows that even with DAF integration, EdgeOL still consistently outperforms all these methods. As shown in the table, EdgeOL provides 2.1$\times$, 2.2$\times$, 2.6$\times$, and 2.2$\times$ energy savings, respectively, while delivering 1.76%, 2.10%, 2.17%, and 1.89% higher accuracy.
EdgeOL outperforms Egeria and AutoFreeze due to its more flexible and finer-grained layer-freezing approach. Spe- cifically, EdgeOL assesses layers individually rather than in modules (i.e., layer blocks), and it freezes all identified converged layers without forcing layers to be frozen sequentially from front to back. Hence, it avoids overtraining already converged layers in the middle of a non-converged module or after a non-converged layer. In contrast to RigL, EdgeOL effectively addresses sparse training challenges, such as GPU underutilization and workload imbalance. Compared to Ekya, EdgeOL eliminates the inefficiency of Ekya’s trial-and-error method in training configuration (e.g., which layers to freeze), ensuring more effective and efficient performance improvements.
Table 6. Comparison with SOTA efficient learning methods.
Model
Method
ImgNet to CF10 ImgNet to ImgNet-C
Acc.
Time
Energy
Acc.
Time
Energy
Res50
DAF (base)
85.69
4,975
16.90
72.82
10,012
36.79
Egeria (wang2023egeria, )
85.86
4,450
13.57
71.57
8,660
30.68
AutoFreeze (liu2021autofreeze, )
86.76
4,634
14.69
72.71
9,053
30.90
RigL (evci2020rigging, )
85.53
4,283
13.06
72.31
8,544
30.91
Ekya (bhardwaj2022ekya, )
86.07
4,536
13.74
72.00
8,796
30.60
EdgeOL 88.02 3,927 11.91 74.69 6,604 21.65
MBV2
DAF (base)
83.62
1,566
5.24
68.21
3,677
11.04
Egeria (wang2023egeria, )
84.61
1,387
4.17
68.39
2,784
9.14
AutoFreeze (liu2021autofreeze, )
83.03
1,395
4.20
67.35
3,094
9.39
RigL (evci2020rigging, )
83.67
1,366
4.75
67.87
2,696
9.34
Ekya (bhardwaj2022ekya, )
83.41
1,372
4.12
69.33
2,923
9.78
EdgeOL 85.51 1,286 3.53 70.23 2,425 6.43
DeiT
DAF (base)
86.29
5,232
19.59
70.03
11,650
43.47
Egeria (wang2023egeria, )
86.83
4,276
16.88
70.18
8,867
34.71
AutoFreeze (liu2021autofreeze, )
86.14
4,342
16.35
69.55
8,438
33.66
RigL (evci2020rigging, )
85.90
4,413
17.57
69.73
9,096
36.51
Ekya (bhardwaj2022ekya, )
86.22
4,130
16.16
69.67
8,692
34.27
EdgeOL 87.79 3,461 13.56 71.77 6,877 25.29
5.3. Sensitivity Analysis
Number of inference requests. Figure 17 shows the average inference accuracy and energy consumption under different numbers of inference requests. Note that, all the inference requests are arriving following a Poisson distribution (mattson2020mlperf, ). All results in this section are obtained by fine-tuning ResNet50 on “ImgNet to CF10”. EdgeOL consistently achieves higher accuracy than $Immed.$, while consuming significantly less energy. The figure also reveals that the energy saving offered by EdgeOL increases as the total number of inference requests decreases. This occurs because when the number of inference requests decreases, EdgeOL (achieved by DAF) will decrease fine-tuning frequency, thereby reducing the energy from execution overheads such as system initialization, as explained in Section 3.2.
Data Arrival distribution. In addition to the Poisson distribution, we also evaluate EdgeOL under different arrival distributions for both training data and inference requests, including the uniform distribution (kuipers2012uniform, ), normal distribution (altman1995statistics, ), and a real-world trace from Video Timeline Tags dataset (huang2020multimodal, ). As depicted in Figure 17, EdgeOL consistently excels in both accuracy and energy consumption compared to $Immed$., showing that EdgeOL is adept at handling a variety of scenarios with different data arrival distributions.
CKA variation threshold. In our experiments, a layer whose CKA variation is less than 1% is considered converged (as mentioned in Section 4.2). Additionally, we further evaluate the performance of EdgeOL under various CKA variation thresholds. Figure 17 shows that decreasing the threshold will lead to higher energy consumption and also higher accuracy. However, the accuracy saturates when the threshold is low enough (e.g., 1%).
5.4. Handling Multiple Scenario Changes
We evaluate EdgeOL when there are multiple model deployment scenario changes throughout online learning. We employ two CORe50 benchmarks NICv2_79 and NICv2_391 (lomonaco2020rehearsal, ), which are widely used in testing the model’s adaptability in dynamically changing scenarios. “NIC” stands for New Instance and Classes, and the numbers (79 and 391) indicate the total scenarios in each benchmark. The scenarios appear one after one, each of which introduces both i) new classes of data and ii) instances of existing classes but with new patterns (e.g., different environmental conditions such as changes in illumination and background).
In Table 7, we compare EdgeOL with layer freezing method Egeria (wang2023egeria, ) and efficient online learning framework Ekya (bhardwaj2022ekya, ). Following the experimental setup described in Section 5.2, both Egeria and Ekya utilize our inter-tuning level optimization, DAF, to ensure a fair comparison. EdgeOL achieves energy savings of 2.3$\times$ and 2.0$\times$ over Egeria and Ekya, respectively, while also providing 1.15% and 1.00% higher accuracy. The superior performance of EdgeOL in handling multiple scenario changes can be attributed to the Reset & Resume design (Section 4.3). When encountering a new scenario, it enables a quick adaptation by increasing the fine-tuning frequency and selectively resuming fine-tuning only for those frozen layers that become unstable in the new scenario.
Table 7. Experiments on handling multiple scenario changes.
Model
Method
NICv2_79 NICv2_391
Acc.
Time
Energy
Acc.
Time
Energy
Res50
DAF (base)
56.94
9,718
34.07
54.20
10,113
34.40
Egeria (wang2023egeria, )
56.60
8,634
29.19
54.23
9,334
29.62
Ekya (bhardwaj2022ekya, )
57.01
8,586
28.80
54.54
9,281
29.00
EdgeOL 57.86 7,482 23.50 55.73 7,720 24.08
MBV2
DAF (base)
53.28
3,066
10.00
50.87
3,224
10.21
Egeria (wang2023egeria, )
53.60
2,796
8.55
50.95
2,916
8.56
Ekya (bhardwaj2022ekya, )
52.78
2,759
8.31
51.22
2,903
8.37
EdgeOL 54.11 2,483 6.70 51.94 2,633 6.81
DeiT
DAF (base)
56.42
9,845
37.19
53.63
10,583
39.91
Egeria (wang2023egeria, )
56.48
8,302
33.90
53.11
8,040
33.02
Ekya (bhardwaj2022ekya, )
56.73
7,885
31.77
53.56
8,844
34.18
EdgeOL 57.51 6,490 25.66 54.70 7,090 28.33
Table 8. Experimental results in semi-supervised learning. The test case is “ImgNet to CF10”.
Model
Method
Acc.
Time (Sec.)
Energy (Wh)
ResNet50
Immed.
63.10
35,757
84.60
EdgeOL 64.38 6,142 25.40
MobileNetV2
Immed.
57.88
5,628
13.92
EdgeOL 59.07 1,761 7.01
DeiT
Immed.
63.30
17,460
40.48
EdgeOL 64.52 4,163 15.33
5.5. Semi-supervised Learning
Next, we evaluate the ability of our EdgeOL to utilize unlabeled data by applying semi-supervised learning. We choose the common configuration that only 10% of the training data is labeled (zhai2019s4l, ; wang2022np, ). As shown in Table 8, compared to $Immed.$, EdgeOL delivers 1.23% higher accuracy and saves 76% of execution time and 61% of energy, on average. These results demonstrate that EdgeOL works well in semi-supervised learning scenarios. This is because both DAF and SimFreeze are robust to the insufficient labeled data as i) DAF only needs a very small amount of labeled validation data to get the validation accuracy to adjust the fine-tuning frequency and ii) SimFreeze freezes layers by self-representational similarity, which can be acquired without data labels.
Table 9. Average inference accuracy when quantization is applied. The results are obtained on ResNet50.
Method
ImgNet to CF10 ImgNet to ImgNet-C
8-bit
32-bit
8-bit
32-bit
Immed.
85.25
85.86
72.38
72.95
EdgeOL
87.93
88.02
74.07
74.69
5.6. Compatibility with Quantization
We also evaluate the compatibility of our EdgeOL with quantization technique (stock2020training, ; gupta2015deep, ). We apply 8-bit fixed-point quantization to weights, activations, the gradient of weights, and the gradient of activations. Following the prior works, we compare the accuracy results since the simulated quantization-aware training is used (zhou2016dorefa, ; wang2018training, ; zhu2020towards, ; zhao2021distribution, ). Table 9 shows that EdgeOL outperforms immediate online learning in 32-bit floating-point baselines with an accuracy improvement of 1.95%. Additionally, when employing 8-bit quantization, EdgeOL achieves a 2.19% higher accuracy. These results suggest that EdgeOL’s advantages are maintained when quantization techniques are used, demonstrating compatibility and robustness.
6. Related Work
A number of approaches have been proposed to reduce the computation costs of DNN models, thereby reducing energy and execution time. E2Train (wang2019e2, ) proposes to drop mini-batches randomly, skip layers selectively, and use low-precision back-propagation during training to reduce the computation costs. (DBLP:conf/iclr/YouL0FWCBWL20, ) designs a low-cost method to train the small but critical subnetworks to achieve the same accuracy as the original neural networks. (hu2022lora, ) proposes to use lightweight low-rank matrices to adapt the weights of original models, slightly sacrificing model representational power to reduce the training costs. However, these and most other prior works focus on offline learning. For online learning, some methods are proposed to filter important data for training to minimize the cost (panda2016conditional, ; wu2021enabling, ). These approaches can effectively reduce training costs. Nevertheless, our approach is complementary to them since it focuses on determining the moment to trigger fine-tuning adaptively and freezing layers selectively.
7. Conclusion
In this paper, we design an efficient and accurate online framework for edge devices, namely EdgeOL. It addresses requirements for both adaptiveness and energy efficiency for efficient online learning from both inter- and intra-tuning levels. Our experiments show that EdgeOL significantly reduces training time and energy consumption while simultaneously improving inference accuracy.